OpenStudy (janu16):

What is the equation of the quadratic graph with a focus of (4, 3) and a directix of y = 13?

OpenStudy (janu16):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

as the focus is below the directrix then the graph is concave down. line of symmetry is x=4.  distance between the focus and directrix is 2 multiplied by the focal length. distance is 2b = -10 => focal length is b = -5 units vertex is (4. 3 - (-5)) => (4. 8) (x^2−h)=4a(y−k) (x−4^)2=4*(−5)(y−8) now simplify all this and get your equation

OpenStudy (alexandervonhumboldt2):

woops mistake wait

OpenStudy (alexandervonhumboldt2):

this is correct distance between the focus and directrix is 2 multiplied by the focal length. distance is 2b = -10 => focal length is b = -5 units vertex is (4. 3 - (-5)) => (4. 8) (x^2−h)=4b(y−k) (x−4)^2=4*(−5)(y−8) now simplify all this and get your equation

OpenStudy (janu16):

so do you simplify both of it?

OpenStudy (alexandervonhumboldt2):

not both you simplify this equation (x−4)^2=4*(−5)(y−8)

OpenStudy (janu16):

= −one twentieth (x − 4)2 + 8

OpenStudy (janu16):

?